OpenStudy (anonymous):

f(x) = x^2 + 6 Find using the difference quotient f(a + h) − f(a)/h

OpenStudy (agent0smith):

This will help you get started.... f(a+h) means you need to plug in a+h in place of x....\[\Large f(a+h) = (x+h)^2 +6\]^ now foil/simplify that. same thing for f(a)\[\Large f(a) = a^2 + 6\] Then you want to plug those into (f(a + h) − f(a))/h

OpenStudy (anonymous):

First find f(a+ h). To find this you need to substitute (a + h) in for x f(a+h) = ( a+h) ^2 + 6 f(a+h) = (a+h)(a+h) +6 f(a + h )= a^2 + 2ah + h^2 + 6 Now find f(a) f(a) = x^2 +6 f(a) = a^2 + 6 Now substitute all of this in ! Literally do a cut and paste ! Where you have f(a+h), paste in a^2 +2ah +h^2 +6 Now, you must also paste in what you found for f(a) ! When you paste this in, you need to put it in parentheses preceded by a negative! We must do f(a+h) - f(a) So, the numerator would look like a: (a^2 +2ah +h^2+6) -(a^2+6) This simplifies to a^2 +2ah +h^2 +6 -a^2 -6 2ah +h^2 Now, all of this must be put over h Factor h out of the numerator h(2a +h) / h Cancel the h out and you are left with 2a+h

OpenStudy (anonymous):

thank you!!

OpenStudy (anonymous):

f(x) = 2/x + 5 FInd using difference quotient

OpenStudy (anonymous):

f(a)=2/a+5 f(a+h)= 2/(a+h)+5

OpenStudy (agent0smith):

You should post it as a new question. But you'll just have to add the fractions in that one too.